The graphs show the demand for two goods by a single individual. 1. The price for good A and B start out at $20 and the individuals demand is D2 on both graphs 2. Then the price of good A rises to $70. Use the endpoint method for all elasticity calculations Good A Good B 3140 140 Q 140
media%2Fe03%2Fe036c0e8-b399-4a98-a9ae-4e
Answer #1

Cross price Ed = % change in quantity demanded of Good B/% change in price of Good A

P1 = 20, P2 = 70, Q1 = 105, Q2 = 140

Cross price Ed = (P1 + P2)/(Q1 + Q2) x (Q2 - Q1)/(P2 - P1)

= 90/245 x 35/50 = 63/245 = 0.26